OpenStudy (anonymous):

Help Please! Reduce rational expression, 4x+16 x-4 ______________________ . ______________________ x to the 2 power -16 x Thank you.

OpenStudy (anonymous):

(4x+16)/(x^2-16) * (x-4)/x ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

You need to simplify each term first. What can 4x+16 be written as? Similarly, what can x^2-16 be written as? Then, cancel out same terms when they occur in numerator and denominator.

OpenStudy (anonymous):

Factor out 4 from 4x+16= 4(x+4) For x^2-16 use difference of squares =(x-4)(x+4) and follow GT instructions

OpenStudy (anonymous):

i got x+4 over x... or is it x+4 over 4x

OpenStudy (anonymous):

It should be 4/x.

OpenStudy (anonymous):

4(x+4) * (x-4) / (x+4)(x-4)x = 4/x

OpenStudy (anonymous):

Thank you